# SPIN-Bench: How Well Do LLMs Plan Strategically and Reason Socially?

### v1

@@ 1 Introduction
Beyond step-wise decision-making, modern AI systems increasingly need social intelligence: the ability to negotiate, cooperate, and reason about other agents’ hidden goals and beliefs (Hou et al., 2024; Cross et al., 2024). Environments such as Diplomacy († et al.(2022)(FAIR)†, Bakhtin, Brown, Dinan, Farina, Flaherty, Fried, Goff, Gray, Hu, et al., FAIR), Avalon (Light et al., 2023), Werewolf (Xu et al., 2023), and various negotiation-based games (Liang et al., 2023; Abdelnabi et al., 2023; Hua et al., 2024) require not just logical strategies but also alliance formation, perspective-taking, and the capacity to handle incomplete or uncertain information.
@@ 2.1 Evaluation of LLM in Planning
LLMs have shown both promise and notable limitations in tasks demanding methodical, step-wise planning. Early work highlighted fundamental shortcomings in LLMs’ reasoning abilities, particularly on problem formulations grounded in PDDL (Hao et al., 2023; Valmeekam et al., 2023, 2024a; Wang et al., 2024). Yet, improvements in test-time scaling and prompt engineering have yielded models that can now achieve over 95% accuracy on discrete, place-based PDDL benchmarks (Valmeekam et al., 2024b; Wang et al., 2024). While these developments indicate an upward trajectory for LLM-based planners, current assessments tend to focus on relatively narrow or routine domains (e.g., classical planning formulations or daily schedule generation) (Zheng et al., 2024; Xie et al., 2024). Such specialized settings fail to expose many of the deeper challenges involved in long-horizon and multi-agent scenarios.
To advance LLMs beyond these simpler planning tests, there is a need for more holistic benchmarks that illuminate how models handle large state and action spaces, manage partial or uncertain information, and adapt their strategies to non-trivial dynamics. As we argue in this paper, such a ”north star” benchmark should not only evaluate how LLMs perform step-wise or fact-based reasoning, but also capture the interplay of conceptual inference (e.g., hidden intentions, shifting goals) and social intelligence (e.g., negotiation, alliance formation). These dimensions are especially vital if LLMs are to fulfill roles in complex decision-support systems and human–AI teaming.
2.2 LLMs for Gaming
Games—whether board, card, or digital—offer an appealing testbed for LLMs because they often demand multi-step reasoning, strategic planning, and, in many instances, multi-agent interaction. Recent work on single-game evaluations has explored domains such as Minecraft (Gong et al., 2024), Avalon (Light et al., 2023), and Werewolf (Xu et al., 2023). While these studies provide insights into how LLMs perform under specific game mechanics, each game’s unique rules and constraints limit how much we can generalize about broader AI competencies.
Efforts to diversify gaming benchmarks have led to platforms that incorporate multiple tasks and game types (Zhang et al., 2024; Duan et al., 2024b, a; Costarelli et al., 2024), as well as multi-agent frameworks emphasizing coordination or competition (Huang et al., 2024; Cross et al., 2024; Agashe et al., 2024). Although these works have substantially enhanced our understanding of LLMs’ strategic reasoning, many still focus on narrow slices of game-play—for example, analyzing deception or coordination in isolation, using coarse-grained outcome metrics, or constraining interaction to simplified environments. Some critical dimensions—like open-ended negotiation or the evolving interplay between cooperation and conflict—remain yet underexplored.
Other lines of research delve deeper into social intelligence in game-based contexts, investigating theory of mind (Cross et al., 2024; Hou et al., 2024), game-theoretic decision-making (Chen et al., 2023a; Akata et al., 2023; Abdelnabi et al., 2023), and dialogue-driven challenges (Qiao et al., 2023; Chen et al., 2024; Duan et al., 2024b; Abdulhai et al., 2023). However, many existing setups still employ relatively constrained strategies or limited forms of social dynamics, risking an incomplete view of how LLMs behave under truly complex, multi-turn interactions.
In contrast, our SPIN-Bench provides a unified setting that encompasses formal PDDL analysis, multi-agent cooperation and competition, open-ended dialogues, and direct comparisons against optimal solvers and human baselines. By weaving strategic planning and social intelligence assessments into a single testbed, we aim to uncover richer insights into the interplay between step-wise action selection and conceptual inference in extended-horizon, multi-agent gaming environments.
@@ 3.1 Problem Definition: Three Settings for Action Sequence Generation
Strategic games extend multi-agent scenarios by introducing dynamic alliances, partial cooperation, and negotiation. Here, the environment state may include past communication logs or partially hidden objectives, allowing agents to form or break alliances, re-evaluate shared goals, and even negotiate mid-game. This richer structure captures a broader spectrum of realistic multi-agent behavior that blends both cooperation and competition within the same session—Diplomacy († et al.(2022)(FAIR)†, Bakhtin, Brown, Dinan, Farina, Flaherty, Fried, Goff, Gray, Hu, et al., FAIR) being a prominent example.
@@ 3.2 Game Taxonomy and Environments
Within SPIN-Bench, classical planning tasks offer a controlled environment to rigorously test core reasoning skills. We employ three subcategories: (i) factual retrieval in domains like elevator, grid, and floortile with progressively expanding state spaces; (ii) spatial reasoning in a modified floortile scenario that challenges models to track agent location from relative instructions; (iii) full classical planning tasks spread across 21 domains (1,280 tasks), each involving sequential decisions, resource management, and constraint handling. Appendix A.1 provides domain-specific details. By starting with simpler tasks and escalating to complex multi-step planning, SPIN-Bench pinpoints whether an LLM’s failures arise from inadequate state tracking, partial-order reasoning, or chain-of-thought breakdowns.
@@
To evaluate adversarial reasoning, we include three widely studied turn-based board games of escalating complexity: Tic-tac-toe, Connect Four, and Chess. Each underscores a key facet of competitive strategy: from short-range lookahead and forcing moves to deeper multi-step tactics. These games also enable direct comparisons against established AI solvers (Minimax (von Neumann, 1928) or Stockfish (Developers, 2023)) and against human-level baselines, providing a clear gauge of how far LLMs are from optimal or near-optimal play.
@@
For multi-agent collaboration under imperfect information, SPIN-Bench leverages the cooperative card game Hanabi. Here, each player observes others’ cards but not their own, and must communicate implicitly to play cards in the correct order. This format demands not just straightforward planning but also trust-building, inference about hidden states, and coordinated actions — essential for more advanced cooperative planning.
@@
Finally, SPIN-Bench incorporates Diplomacy, a complex multi-player board game in which negotiation, alliance formation, and strategic betrayal are integral. Players exchange messages to coordinate or mislead, while simultaneously issuing movement orders to capture territories. Because negotiation is central, LLMs are tested on both long-horizon planning and social intelligence—managing alliances, inferring hidden intentions, and adapting their strategies in response to evolving alliances. This setting is especially suited for evaluating how chain-of-thought reasoning coexists with dynamic social interaction under partial cooperation and partial competition.
@@ 3.3 Benchmark Construction
To accommodate a broad range of tasks (from single-agent PDDL to multi-agent diplomacy), we implement a flexible interface that feeds LLMs the current state description, relevant history, and a list of legal actions or moves. In dialogue-based games (like Diplomacy), we also provide any public or private messages. Models are allowed up to ten retries when proposing illegal moves; exceeding this limit results in an automatic loss. This setup strikes a balance between strict rule compliance and capturing potential misjudgments or misunderstandings that might arise in open-ended text generation.
@@ 3.4 Evaluation Metrics
where is the total number of tasks, indicates correctness (0 or 1), and is the required number of steps for task . This weighted measure highlights performance gaps on more complex problems. In competitive games, we compare LLM moves against solver-generated top- actions and track head-to-head outcomes against the solver. For Hanabi, we use the final score to assess cooperative efficiency. Meanwhile, Diplomacy performance is evaluated through factual consistency, order-level correctness, and final game outcomes (see Appendix A.4 for details).
@@
We maintain internal Elo ratings for competitive games, updating a model’s rating after each match to reflect its performance relative to an opponent’s rating. This helps us gauge the ”skill gap” among different LLMs and compare them against human baselines (Appendix B.1).
@@
In negotiation-heavy settings like Diplomacy, we use six fine-grained, LLM-assisted metrics: (1) alignment of messages with stated negotiation strategies, (2) proposal acceptance rate, (3) mutual vs. one-sided benefit of proposed deals, (4) peace vs. conflict inclination, (5) evidence of perspective-taking, and (6) conditional negotiation.
An advanced reference model (e.g., o1) is prompted to annotate chat logs along these dimensions, enabling a deeper look into each LLM’s social intelligence, strategic coherence, and ability to balance persuasion with deception. Details of these annotations appear in Appendix A.4.2.
@@ 4 Experiments on SPIN-Bench
| Classical Planning | Competitive Games | Collaborative: Hanabi | |||||||||
| Model | Plan Acc | N-Step | TTT | C4 | CH | C4 | CH | 2P | 3P | 4P | 5P |
| o1 | 58.59 | 16.09 | 30.0 | 100.0 | 100.0 | 83.1 | 45.9 | 16.4 | 14.8 | 14.8 | 14.2 |
| o1-mini | 13.20 | 1.95 | 50.0 | 100.0 | 100.0 | 87.0 | 36.5 | 6.8 | 7.4 | 11.4 | 10.2 |
| o3-mini | 51.25 | 13.04 | 80.0 | 100.0 | 100.0 | 74.2 | 52.8 | 8.8 | 7.6 | 8.8 | 8.0 |
| GPT-4o | 8.75 | 0.60 | 100.0 | 100.0 | 100.0 | 84.1 | 32.2 | 6.6 | 4.8 | 4.8 | 4.6 |
| GPT-4-turbo | 5.62 | 0.13 | 40.0 | 100.0 | 100.0 | 83.8 | 38.7 | 5.2 | 5.6 | 5.0 | 6.0 |
| Claude 3.5 Sonnet | 20.55 | 4.44 | 40.0 | 100.0 | 100.0 | 78.9 | 49.5 | 8.2 | 9.4 | 7.4 | 8.4 |
| Claude 3.5 haiku | 4.22 | 0.30 | 50.0 | 100.0 | 100.0 | 69.6 | 35.9 | 2.4 | 4.0 | 2.8 | 2.8 |
| DeepSeek R1 | 44.30 | 10.71 | 90.0 | 100.0 | 100.0 | 78.9 | 47.8 | 6.0 | 16.0 | 11.3 | 13.0 |
| Llama-3.3-70b | 5.78 | 0.32 | 100.0 | 100.0 | 100.0 | 79.5 | 25.4 | N/A | N/A | N/A | N/A |
| 4 Agents (w/o vs. w. Negotiation) | Negotiation Metrics: Social Intelligence | ||||||||||||
| Model | BS | Move | Attack | SS | SO | SC | CR | AlignR | AcceptR | Mut:1S | P:C | Persp. | Cond. |
| o1 | pass | 0.70/0.85 | 0.44/0.83 | 0.45/0.46 | 0.00/0.00 | 17/10 | 37/18 | 0.9800 | 0.6533 | 20.50 | 7.05 | 0.333 | 0.196 |
| GPT-4o | pass | 0.55/0.62 | 0.33/0.37 | 0.30/0.29 | 0.25/0.40 | 15/17 | 28/29 | 0.9545 | 0.6238 | 31.50 | 11.00 | 0.310 | 0.038 |
| GPT-4-turbo | fail | 0.47/0.66 | 0.14/0.36 | 0.23/0.30 | 0.00/0.43 | 7/8 | 14/19 | 0.9643 | 0.5411 | 52.00 | 2.49 | 0.416 | 0.077 |
| DeepSeek-R1 | fail | N/A | N/A | N/A | N/A | N/A | N/A | 0.9571 | 0.6071 | 3.38 | 0.77 | 0.322 | 0.406 |
| o1-preview | pass | N/A | N/A | N/A | N/A | N/A | N/A | 0.9490 | 0.4699 | 5.00 | 1.60 | 0.400 | 0.173 |
| Claude 3.5 Haiku | fail | N/A | N/A | N/A | N/A | N/A | N/A | 0.9067 | 0.3571 | 8.75 | 0.84 | 0.225 | 0.083 |
@@
In what follows, we first outline the overall experimental setup (Section 4.1) and a glance over the model performance (Section 4.2). Next, we detail results on core planning skills (Section 4.3), examine the impact of action complexity and agent scaling (Section 4.4), evaluate social intelligence and negotiation behaviors (Section 4.5), and conclude with a summary of key insights (Section 4.6).

### v3

@@ 1 Introduction
Beyond step-wise decision-making, modern AI systems increasingly need social intelligence: the ability to negotiate, cooperate, and reason about other agents’ hidden goals and beliefs (Hou et al., 2024; Cross et al., 2024). Environments such as Diplomacy (Meta Fundamental AI Research Diplomacy Team et al.(2022)Meta Fundamental AI Research Diplomacy Team (FAIR), Bakhtin, Brown, Dinan, Farina, Flaherty, Fried, Goff, Gray, Hu, et al., FAIR), Avalon (Light et al., 2023), Werewolf (Xu et al., 2023), and various negotiation-based games (Liang et al., 2023; Abdelnabi et al., 2023; Hua et al., 2024) require not just logical strategies but also alliance formation, perspective-taking, and the capacity to handle incomplete or uncertain information.
@@ 2.1 Evaluation of LLM in Planning
LLMs have shown both promise and notable limitations in tasks demanding methodical, step-wise planning. Early work highlighted fundamental shortcomings in LLMs’ reasoning abilities, particularly on problem formulations grounded in PDDL (Hao et al., 2023; Valmeekam et al., 2023, 2024a; Wang et al., 2024). Yet, improvements in test-time scaling and prompt engineering have yielded models that can now achieve over 95% accuracy on discrete, place-based PDDL benchmarks (Valmeekam et al., 2024b; Wang et al., 2024). While these developments indicate an upward trajectory for LLM-based planners, current assessments tend to focus on relatively narrow or routine domains (e.g., classical planning formulations or daily schedule generation) (Zheng et al., 2024; Xie et al., 2024). Such specialized settings fail to expose many of the deeper challenges involved in long-horizon and multi-agent scenarios. To advance LLMs beyond these simpler planning tests, there is a need for more holistic benchmarks that illuminate how models handle large state and action spaces, manage partial or uncertain information, and adapt their strategies to non-trivial dynamics. These dimensions are especially vital if LLMs are to fulfill roles in complex decision-support systems and human–AI teaming.
LLMs for gaming Games offer an appealing testbed for LLMs because they often demand multi-step reasoning, strategic planning, and multi-agent interaction. Recent work on single-game evaluations has explored domains such as Minecraft (Gong et al., 2024), Avalon (Light et al., 2023), and Werewolf (Xu et al., 2023), but these narrow settings limit generalizability. Broader benchmarks incorporate multiple game types (Paglieri et al., 2024; Ruoss et al., 2024; Zhang et al., 2024; Duan et al., 2024b, a; Costarelli et al., 2024), and multi-agent frameworks emphasizing coordination or competition (Huang et al., 2024; Cross et al., 2024; Agashe et al., 2024), yet critical dimensions—like open-ended negotiation, evolving cooperation versus conflict, and richer social dynamics—remain underexplored. Other lines of research delve deeper into social intelligence in game-based contexts, investigating theory of mind (Cross et al., 2024; Hou et al., 2024), game-theoretic decision-making (Chen et al., 2023a; Akata et al., 2023; Abdelnabi et al., 2023), and dialogue-driven challenges (Qiao et al., 2023; Chen et al., 2024; Duan et al., 2024b; Abdulhai et al., 2023), but often within constrained strategies or limited forms of social dynamics. In contrast, our SPIN-Bench unifies strategic planning and social intelligence in a single testbed - combining formal PDDL analysis, multi-agent cooperation and competition, open-ended dialogues, and direct comparisons against optimal solvers and human baselines-to uncover richer insights into the interplay between step-wise action selection and conceptual inference in extended-horizon gaming environments.
@@ 3.1 Problem Definition: Three Settings for Action Sequence Generation
Strategic games extend multi-agent scenarios by introducing dynamic alliances, partial cooperation, and negotiation. Here, the environment state may include past communication logs or partially hidden objectives, allowing agents to form or break alliances, re-evaluate shared goals, and even negotiate mid-game. This richer structure captures a broader spectrum of realistic multi-agent behavior that blends both cooperation and competition within the same session—Diplomacy (Meta Fundamental AI Research Diplomacy Team et al.(2022)Meta Fundamental AI Research Diplomacy Team (FAIR), Bakhtin, Brown, Dinan, Farina, Flaherty, Fried, Goff, Gray, Hu, et al., FAIR) being a prominent example.
@@ 3.2 Game Taxonomy and Environments
Within SPIN-Bench, classical planning tasks offer a controlled environment to rigorously test core reasoning skills. We employ three subcategories: (i) factual retrieval in domains like elevator, grid, and floortile with progressively expanding state spaces; (ii) spatial reasoning in a modified floortile scenario that challenges models to track agent location from relative instructions; (iii) full classical planning tasks spread across 21 domains (1,280 tasks), each involving sequential decisions, resource management, and constraint handling. Appendix B.1 provides domain-specific details. By starting with simpler tasks and escalating to complex multi-step planning, SPIN-Bench pinpoints whether an LLM’s failures arise from inadequate state tracking, partial-order reasoning, or chain-of-thought breakdowns.
@@
To evaluate adversarial reasoning, we include three widely studied turn-based board games of escalating complexity: Tic-tac-toe, Connect Four, and Chess. Each underscores a key facet of competitive strategy: from short-range lookahead and forcing moves to deeper multi-step tactics. Game introduction is shown in Appendix B.2. These games also enable direct comparisons against established AI solvers (Minimax (von Neumann, 1928) or Stockfish (The Stockfish developers, 2025)) and against human-level baselines, providing a clear gauge of how far LLMs are from optimal or near-optimal play.
@@
For multi-agent collaboration under imperfect information, SPIN-Bench leverages the cooperative card game Hanabi. Here, each player observes others’ cards but not their own, and must communicate implicitly to play cards in the correct order. This format demands not just straightforward planning but also trust-building, inference about hidden states, and coordinated actions — essential for more advanced cooperative planning. Game introduction is shown in Appendix B.3.
@@
Finally, SPIN-Bench incorporates Diplomacy, a complex multi-player board game in which negotiation, alliance formation, and strategic betrayal are integral. Players exchange messages to coordinate or mislead, while simultaneously issuing movement orders to capture territories. Game introduction and tutorial are shown in Appendix B.4. Because negotiation is central, LLMs are tested on both long-horizon planning and social intelligence—managing alliances, inferring hidden intentions, and adapting their strategies in response to evolving alliances. This setting is especially suited for evaluating how chain-of-thought reasoning coexists with dynamic social interaction under partial cooperation and partial competition. It captures a broader spectrum of realistic multi-agent behavior that blends both cooperation and competition within the same session.
@@ 3.3 Benchmark Construction
To accommodate a broad range of tasks (from single-agent PDDL to multi-agent diplomacy), we implement a flexible interface that feeds LLMs the current state description, relevant history, and a list of legal actions or moves. We also design game-specific Chain of Thought(CoT) (Wei et al., 2022) prompt for the game agent. In dialogue-based game Diplomacy, we use the Diplomacy game engine (Paquette et al., 2020) and also provide any public or private messages. Models are allowed up to ten retries when proposing illegal moves; exceeding this limit results in an automatic loss. This setup strikes a balance between strict rule compliance and capturing potential misjudgments or misunderstandings that might arise in open-ended text generation. Full details on engine integration, prompt templates and domain rules, appear in Appendix B, D.
@@ 3.4 Evaluation Metrics
where is the total number of tasks, indicates correctness (0 or 1), and is the required number of steps for task . This weighted measure highlights performance gaps on more complex problems. In competitive games, we compare LLM moves against solver-generated top- actions and track head-to-head outcomes against the solver. For Hanabi, we use the final score to assess cooperative efficiency. Meanwhile, Diplomacy performance is evaluated through factual consistency, order-level correctness, and final game outcomes (see Appendix B.4 for details).
@@
We maintain internal Elo ratings (do not directly reflect performance among humans) for competitive games, updating a model’s rating after each match to reflect its performance relative to an opponent’s rating. This helps us gauge the ”skill gap” among different LLMs and compare them against human baselines.
@@
In negotiation-heavy settings like Diplomacy, we use six fine-grained, LLM-assisted metrics: (1) alignment of messages with stated negotiation strategies, (2) proposal acceptance rate, (3) mutual vs. one-sided benefit of proposed deals, (4) peace vs. conflict inclination, (5) evidence of perspective-taking, and (6) conditional negotiation. We then prompt Claude 3.7 Sonnet to annotate chat logs along these dimensions, enabling a deeper look into each LLM’s social intelligence, strategic coherence, and ability to balance persuasion with deception. Metric details appear in Appendix B.4.2. We also conduct human annotation experiments to verify the accuracy of LLM annotations in Appendix C.2.3.
@@ 4 Experiments on SPIN-Bench
| Classical Planning | Competitive Games | Collaborative: Hanabi | Avg. | |||||||||
| Model | Plan Acc | N-Step | TTT | C4 | CH | C4 | CH | 2P | 3P | 4P | 5P | Score |
| o1 | 58.59 | 16.09 | 70.0 | 0.0 | 0.0 | 83.1 | 45.9 | 16.4 | 14.8 | 14.8 | 14.2 | 49.8 |
| o1-mini | 13.20 | 1.95 | 50.0 | 0.0 | 0.0 | 87.0 | 36.5 | 6.8 | 7.4 | 11.4 | 10.2 | 33.0 |
| o3-mini | 51.25 | 13.04 | 20.0 | 0.0 | 0.0 | 74.2 | 52.8 | 8.8 | 7.6 | 8.8 | 8.0 | 33.1 |
| GPT-4o | 8.75 | 0.60 | 0.0 | 0.0 | 0.0 | 84.1 | 32.2 | 6.6 | 4.8 | 4.8 | 4.6 | 20.8 |
| GPT-4-turbo | 5.62 | 0.13 | 60.0 | 0.0 | 0.0 | 83.8 | 38.7 | 5.2 | 5.6 | 5.0 | 6.0 | 27.5 |
| Claude 3.5 Sonnet | 20.55 | 4.44 | 60.0 | 0.0 | 0.0 | 78.9 | 49.5 | 8.2 | 9.4 | 7.4 | 8.4 | 34.3 |
| Claude 3.5 Haiku | 4.22 | 0.30 | 50.0 | 0.0 | 0.0 | 69.6 | 35.9 | 2.4 | 4.0 | 2.8 | 2.8 | 20.8 |
| DeepSeek R1 | 44.30 | 10.71 | 10.0 | 0.0 | 0.0 | 78.9 | 47.8 | 6.0 | 16.0 | 11.3 | 13.0 | 36.6 |
| Llama-3.3-70b | 5.78 | 0.32 | 0.0 | 0.0 | 0.0 | 79.5 | 25.4 | 2.4 | 0.8 | 0.6 | 1.4 | 13.1 |
| 4 Agents (w/o vs. w. Negotiation) | Negotiation Metrics: Social Intelligence | |||||||||||
| Model | Move | Attack | SS | SO | SC | CR | AlignR | AcceptR | Mut:1S | P:C | Persp. | Cond. |
| o1 | 0.70/0.85 | 0.44/0.83 | 0.45/0.46 | 0.00/0.00 | 17/10 | 37/18 | 1.000 | 0.673 | 100.0 | 9.26 | 0.33 | 0.36 |
| GPT-4o | 0.55/0.62 | 0.33/0.37 | 0.30/0.29 | 0.25/0.40 | 15/17 | 28/29 | 0.994 | 0.373 | 36.6 | 3.16 | 0.33 | 0.04 |
| GPT-4-turbo | 0.47/0.66 | 0.14/0.36 | 0.23/0.30 | 0.00/0.43 | 7/8 | 14/19 | 0.995 | 0.437 | 166.0 | 2.76 | 0.45 | 0.03 |
| DeepSeek-R1 | 0.64/0.65 | 0.47/0.47 | 0.35/0.43 | 0.00/0.67 | 16/16 | 22/27 | 0.982 | 0.416 | 5.6 | 0.95 | 0.21 | 0.51 |
| o1-preview | 0.82/0.87 | 0.69/0.67 | 0.46/0.40 | 0.00/0.17 | 15/11 | 38/26 | 0.994 | 0.479 | 27.3 | 2.04 | 0.44 | 0.15 |
| Claude 3.5 Haiku | 0.48/0.37 | 0.23/0.20 | 0.38/0.07 | 0.00/0.00 | 0/5 | 1/7 | 0.996 | 0.409 | 43.0 | 0.83 | 0.47 | 0.07 |
@@
In what follows, we first outline the overall experimental setup (Section 4.1) and a glance over the model performance (Section 4.2). Next, we detail results on core planning skills (Section 4.3), examine the impact of action complexity and agent scaling (Section 4.4), evaluate social intelligence and negotiation behaviors (Section 4.5), and conclude with a summary of key insights (Section 4.6). The reproducibility statement of all our evaluations is in Appendix A. Our trajectory viewer present game trajectories and examples in all benchmark settings.